# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v4

@@ Appendix D Evaluation Metric
  • Fréchet Inception Distance (FID) [31]
@@
    Methodology: Uses the Inception-v3 network [73] to extract features. Assumes both feature distributions are multivariate Gaussian and computes the Fréchet distance between them.
@@
  • Spatial Fréchet Inception Distance (sFID) [59]
@@
  • Precision and Recall for Distributions (Precision/Recall) [44]
@@ Appendix E Methods for Comparison
  • ADM [19] enhances U-Net-based architectures for diffusion models and introduces classifier-guided sampling, a technique used to balance the quality-diversity tradeoff and improve overall performance.
@@
  • VDM++ [40] improves diffusion model training efficiency through an adaptive noise scheduling mechanism that dynamically adjusts noise levels during optimization.
@@
  • Simple Diffusion [35] targets high-resolution synthesis by simplifying both network architectures and noise schedules through systematic exploration of lightweight design choices.
@@
  • CDM [33] achieves high-fidelity generation via a cascaded pipeline: training low-resolution diffusion models first, then progressively refining details through super-resolution diffusion stages.
@@
  • LDM [69] accelerates training while maintaining generation quality by operating in compressed latent spaces, modeling image distributions at reduced dimensionality compared to pixel space.
@@
  • DiT [62] employs a pure transformer backbone for diffusion models, incorporating AdaIN-zero modules as an important component modification against vanilla vision transformer [20] of its architecture.
@@
  • SiT [56] provides an extensive analysis of how the training efficiency of DiTs can be improved by transitioning from discrete diffusion to continuous flow-based modeling, which can be seen as a flow-based version of DiT.
@@
  • SD-DiT [86] leverages IBOT’s [85] training paradigm that combines DINO loss [8] and BEIT loss [4] for efficiently training diffusion transformers.
@@
  • MaskDiT [84] proposes an asymmetric encoder-decoder scheme for efficient training of diffusion transformers, where they train the model with an auxiliary mask reconstruction task similar to MAE [29].
@@
  • TREAD [42] introduces a dynamic token routing strategy combined with the mask reconstruction task similar to MAE [29] and MaskDiT [84] to accelerate the training of diffusion models.
@@
  • REPA [81] achieves significant improvements in both training efficiency and generation quality by aligning the latent feature of the diffusion model with that of a large-scale data pre-trained representation model (e.g., DINOv2 [61]).
@@
  • MAETok changes the SD-VAE to MAE-Tok, which is trained with auxiliary mask reconstruction loss and aligns loss with three representation targets (HOG’s [15], DINOv2’s [61], and CLIP’s [65]) and obtains diffusion transformer with better generation performance.
Appendix F Convergence Speed Against REPA
Here we provide a quantitative comparison of convergence speed with SiT-XL baseline and SiT-XL + REPA [81]. It is worth noting that REPA leverages a powerful representation foundation model whose training data and resources are far beyond training a diffusion transformer on ImageNet.
As shown in Figure 8, although REPA converges quickly at the beginning of training due to its large-scale, pre-trained representation model, the performance saturates after about 200 epochs. On the contrary, in SRA, since the teacher’s ability is constantly enhanced during training and thus better and better guidance is provided, continuous performance improvement can be achieved. From another perspective, the representation guidance signal in REPA is fixed in every timestep and epoch, while our supervision signal is dynamically changing. This provides a wider variety of learning cues to help the model continuously improve its performance.
Figure 8: Training epochs vs. FID plot without classifier-free guidance (CFG). The benefit of REPA encounters saturation after approximately 200 epochs, while our method provides continuous performance improvement.
Appendix G Detailed Setup of Ablation Study
@@
Appendix H Ablation Results of DiT
@@
Appendix I More Discussion on Related Work
@@
Diffusion transformers. Currently, the diffusion model is progressively transitioning from a U-Net-based architecture to a Transformer-based one, owing to the superior scalability of the latter. At the beginning, U-ViT [3] shows transformer-based backbones with skip connections can be an effective backbone for training diffusion models. Then, DiT [62] shows skip connections are not even necessary components, and a pure transformer architecture can be a scalable architecture for training denoise-based models. Based on DiT, SiT [56] shows the model can be further improved with continuous stochastic interpolants [2]. Moreover, Stable diffusion 3 [22], Lumina-Image 2.0 [64] and FLUX 1 [45] show pure transformers can be scaled up for challenging text-to-image generation, and this characteristic is also verified by Sora [5], CogvideoX [79] and Wan [76] in text-to-video field. Our work focuses on improving the training of DiT (and SiT) architecture based on our proposed self-representation alignment technique.
@@
Exploring representation capacity of diffusion models. With the success of diffusion models to generate detailed images, many works have attempted to test whether discriminative semantic information can be found in diffusion models. GD [58] and DDAE [77] first observe that the intermediate representations of diffusion models have discriminative properties. Driving from this finding, I-DAE [13] deconstructs diffusion models to be a self-supervised Learner. Moreover, Repfusion [78], DiffusionDet [11], and DreamTeacher [46] use diffusion models to perform various downstream tasks (e.g., semantic segmentation and object detection). Our work also tries to explore the representation capacity of the diffusion model, but we focus on leveraging the representations in diffusion transformers to enhance their generation capacity.
@@
Applying representation guidance to other generation tasks. In addition to pre-training of the class-conditional diffusion model, applying representation guidance can benefit other generation tasks. For example, lbGen [37] utilizes text features from CLIP [65] as a low-biased reference to regulate diffusion model for low-biased dataset synthetics. RCG [47] focuses on unconditional generation, which first use the features from a self-supervised image encoder to train a representation generator and subsequently used it output as the ‘label’ for image generation by a second generator. Diff-AE [63] uses a learnable encoder for discovering the high-level semantics and a diffusion model for modeling stochastic; this dual-encoding improves the realism of the generated images on attribute manipulation and image interpolation tasks. Different from these works, our study focuses on the class-conditional diffusion transformer’s pre-training and exploiting representation guidance in itself and its own training paradigm. But we also believe our plug-and-play method can be easily applied to other tasks with benefits.
@@
Knowledge distillation for diffusion model. Knowledge distillation [27] is also widely used in diffusion models. Its purposes can roughly be divided into two types: improving generation performance and accelerating inference sampling. To achieve the first goal, A more powerful pretrained model is often used to guide the diffusion model during training [16, 81, 24]. For example, TinyFusion [24] combines progressive distillation with architecture-specific optimizations for U-Net backbones, enabling deployment on edge devices. REPA [81] distills the knowledge from a large-scale, pretrained representation foundation model [61, 65] and finds that this distillation can improve both training efficiency and generation quality. Meanwhile, to achieve the second goal, the student and teacher models will be initialized by pre-trained models (here it can be one model or two different models), then the training target is to make the prediction of the student model with fewer steps conform to that of the teacher model with multiple steps [55, 66, 57, 71], thereby achieving the result of accelerated sampling and generation speed. For example, CM [71] constrains the student’s outputs of adjacent points on a sampling path to be consistent with teacher’s. LCM [55] leverages this idea in the latent space. Our study also have some similarities, while we do not need a pretrained model and improve the generation performance of diffusion transformers by applying proposed SRA.
Appendix J More Qualitative Results
@@
Below we show some uncurated generation results on ImageNet 256$\times$256 from the SiT-XL + SRA. We use classifier-free guidance with $w$ = 4.0.

### v5

@@ Appendix D Evaluation Metric
  • Fréchet Inception Distance (FID) [fid]
@@
    Methodology: Uses the Inception-v3 network [inception-model] to extract features. Assumes both feature distributions are multivariate Gaussian and computes the Fréchet distance between them.
@@
  • Spatial Fréchet Inception Distance (sFID) [sfid]
@@
  • Precision and Recall for Distributions (Precision/Recall) [pre-rec]
@@ Appendix E Methods for Comparison
  • ADM [adm] enhances U-Net-based architectures for diffusion models and introduces classifier-guided sampling, a technique used to balance the quality-diversity tradeoff and improve overall performance.
@@
  • VDM++ [vdm] improves diffusion model training efficiency through an adaptive noise scheduling mechanism that dynamically adjusts noise levels during optimization.
@@
  • Simple Diffusion [simplediff] targets high-resolution synthesis by simplifying both network architectures and noise schedules through systematic exploration of lightweight design choices.
@@
  • CDM [cdm] achieves high-fidelity generation via a cascaded pipeline: training low-resolution diffusion models first, then progressively refining details through super-resolution diffusion stages.
@@
  • LDM [ldm] accelerates training while maintaining generation quality by operating in compressed latent spaces, modeling image distributions at reduced dimensionality compared to pixel space.
@@
  • DiT [dit] employs a pure transformer backbone for diffusion models, incorporating AdaIN-zero modules as an important component modification against vanilla vision transformer [vit] of its architecture.
@@
  • SiT [sit] provides an extensive analysis of how the training efficiency of DiTs can be improved by transitioning from discrete diffusion to continuous flow-based modeling, which can be seen as a flow-based version of DiT.
@@
  • SD-DiT [sddit] leverages IBOT’s [ibot] training paradigm that combines DINO loss [dino] and BEIT loss [beit] for efficiently training diffusion transformers.
@@
  • MaskDiT [maskdit] proposes an asymmetric encoder-decoder scheme for efficient training of diffusion transformers, where they train the model with an auxiliary mask reconstruction task similar to MAE [mae].
@@
  • TREAD [tread] introduces a dynamic token routing strategy combined with the mask reconstruction task similar to MAE [mae] and MaskDiT [maskdit] to accelerate the training of diffusion models.
@@
  • REPA [repa] achieves significant improvements in both training efficiency and generation quality by aligning the latent feature of the diffusion model with that of a large-scale data pre-trained representation model (e.g., DINOv2 [dinov2]).
@@
  • MAETok changes the SD-VAE to MAE-Tok, which is trained with auxiliary mask reconstruction loss and aligns loss with three representation targets (HOG’s [hog], DINOv2’s [dinov2], and CLIP’s [clip]) and obtains diffusion transformer with better generation performance.
Appendix F Teacher Network Updating Role
Table 6: Different EMA update rule attempts, other settings are consistent with the default setting of Table 3.2 in the main paper.
[…]
0.0 35.71 42.18
$0.996\xrightarrow{}1.0$ 33.17 44.96
0.9999 29.10 50.20
In other generative learning studies, the EMA model is often used solely for evaluation. However, as we need it to provide guidance during training, we study different updating methods. Here, we investigate three different strategies to build the teacher. First, we find that using teacher copied from a student ($\alpha=0$) would impair the performance, which testifies to our argument in Section 1 . Next, we consider using the strategy widely used in self-supervised learning works [dino, ibot] that updates the momentum coefficient $\alpha$ from 0.996 to 1 during training. However, in our framework, this does not work well. Finally, we use the $\alpha$ of 0.9999 unchanged, which is commonly used in other generative learning works [adm, dit] and find it is also suitable for our framework. In our analysis, first, student copy ($\alpha$= 0.0) performs the worst : this is consistent with the observation in self-supervised learning: student copy leads to unstable training and cannot converge; there is a diffusion loss in our approach, so the training still converges. Second, using $\alpha$= 0.9999 is better than 0.996 → 1.0: In EMA, a larger $\alpha$ indicates a smoother update. We analyze that the reason for keeping such a very large value of $\alpha$ works well in SRA is that: in our method, the enhanced representation learning is ultimately for the service of generation. Thus, when the update of teacher model is relatively intense, the alignment would be less stable, the model may focus more on alignment loss and overlook diffusion loss, thereby destroying some of the original generation behavior. On the contrary, maintaining $\alpha$ in a very large value (0.9999) ensures the alignment target is extremely robust and slow to change, preventing the alignment loss from destabilizing the primary diffusion loss. Thus, we set the momentum coefficient as 0.9999 as our default.
Appendix G Principle of our hyperparameter
Although our method requires some hyperparameter tuning, we believe that some settings are model-independent, while some settings have selection principles. We now give the detailed explanations, and we hope these explanations can help to develop SRA to other new models more easily.
  • Projection Head. In generative models, different layers are often believed to have different generation responsibilities. We hypothesize that using the lightweight projection head instead of directly conducting the alignment can lead to a relatively soft alignment, and could be the reason for avoiding disruption to the model’s original generative behavior. We believe this principle and findings with the design of projection head can be easily transferred to new models.
  • $\alpha$ in EMA decay. In EMA, a larger $\alpha$ indicates a smoother update. We analyze that the reason for keeping such a very large value of $\alpha$ works well in SRA is that: in our method, the enhanced representation learning is ultimately for the service of generation. Thus, when the update of teacher model is relatively intense, the alignment would be less stable, the model may focus more on alignment loss and overlook diffusion loss, thereby destroying some of the original generation behavior. On the contrary, maintaining $\alpha$ in a very large value (0.9999) ensures the alignment target is extremely robust and slow to change, preventing the alignment loss from destabilizing the primary diffusion loss. This can also be indirectly verified by our following analysis of $\lambda$. Given this analysis, we believe the choice of $\alpha$ is will be well-matched to most of the new model.
  • Block layers and time intervals. Principle of the selection of block layers: the block layer for the teacher is selected so that the corresponding representation has strong semantics (See Figure 2 (b)). The block layer for students is the first few layer as the first few layers are more about semantics learning, which is similar to REPA, and discussed in REPA. Principle of the time interval: The intuitive guidance could be that the target representation for the teacher timestep has better semantics, and the teacher timestep is not too far from the current timestep, i.e., a balance between teacher representation semantics and the semantic distance between teacher (timestep sampled from the interval) and student (current timestep). These principles are also applicable to other diffusion transforms (e.g, MMDiT as shown Table 3.3).
[…]
Appendix H Training Speed and GPU Memory Usage Against Baselines
Table 7: Memory useage, per-epoch training time comparison. “Param” counts only the additional parameters required by the auxiliary forward pass. These results are tested with total batch size of 256 on a single machine with 8 A100 GPUs.
Method Mem (GB) Time (h) Param
SiT-B 12.85 0.096
SiT-B + SRA 13.23 0.115 87 M
SiT-L 25.68 0.282
SiT-L + SRA 26.65 0.330 305 M
SiT-XL 29.08 0.394
SiT-XL + SRA 30.24 0.476 481 M
DiT-B 13.96 0.173
DiT-B + SRA 14.78 0.209 87 M
DiT-L 27.85 0.503
DiT-L + SRA 29.26 0.558 305 M
DiT-XL 30.68 0.719
DiT-XL + SRA 32.31 0.804 481 M
As our approach needs one more forward pass on the diffusion backbone to compute the target representation. We now provide a specific comparison on GPU memory usage and training time with baselines. Noting that the extra pass is only for forward, and no gradient computation is needed, so we can use half-precision for fast forward. In addition, it is not needed to pass the whole network, and only a partial pass is needed: 8/12 blocks in SiT-B, 20/28 blocks in SiT-XL. So it can be seen that the additional cost are not very large, and the results in Figure 4 and Table 3.3 also indicate that our method can improve the performance of the model within the same training time.
Appendix I Detailed Setup of Ablation Study
@@
Appendix J Ablation Results of DiT
@@
Appendix K More Discussion on Related Work
@@
EMA model as teacher. Unlike traditional knowledge distillation [kd] that uses a well pre-trained model as the teacher, using EMA model [ema] as the teacher can be seen as self-distillation because the weight of the teacher model is obtained by the weighted moving average of the student. This method often needs a feasible pretext task to succeed. For example, MoCo [moco, mocov3] sets the teacher’s output as a momentum queue and uses contrastive learning to guide the student model; DINO [dino, dinov2] feeds two views of images to the teacher, and the trainable student then forces the student’s output distribution to be close to that of the teacher. Our work also shares some similarities, where we set aligning the student model’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on lower noise of the teacher as our pretext task for training.
@@
Diffusion transformers. Currently, the diffusion model is progressively transitioning from a U-Net-based architecture to a Transformer-based one, owing to the superior scalability of the latter. At the beginning, U-ViT [uvit] shows transformer-based backbones with skip connections can be an effective backbone for training diffusion models. Then, DiT [dit] shows skip connections are not even necessary components, and a pure transformer architecture can be a scalable architecture for training denoise-based models. Based on DiT, SiT [sit] shows the model can be further improved with continuous stochastic interpolants [si]. Moreover, Stable diffusion 3 [sd3], Lumina-Image 2.0 [lumina2] and FLUX 1 [flux] show pure transformers can be scaled up for challenging text-to-image generation, and this characteristic is also verified by Sora [sora], CogvideoX [cogvideox] and Wan [wan] in text-to-video field. Our work focuses on improving the training of DiT (and SiT) architecture based on our proposed self-representation alignment technique.
@@
Exploring representation capacity of diffusion models. With the success of diffusion models to generate detailed images, many works have attempted to test whether discriminative semantic information can be found in diffusion models. GD [gd] and DDAE [ddae] first observe that the intermediate representations of diffusion models have discriminative properties. Driving from this finding, I-DAE [i-dae] deconstructs diffusion models to be a self-supervised Learner. Moreover, Repfusion [repfusion], DiffusionDet [diffusiondet], and DreamTeacher [dreamteacher] use diffusion models to perform various downstream tasks (e.g., semantic segmentation and object detection). Our work also tries to explore the representation capacity of the diffusion model, but we focus on leveraging the representations in diffusion transformers to enhance their generation capacity.
@@
Applying representation guidance to other generation tasks. In addition to pre-training of the class-conditional diffusion model, applying representation guidance can benefit other generation tasks. For example, lbGen [lbgen] utilizes text features from CLIP [clip] as a low-biased reference to regulate diffusion model for low-biased dataset synthetics. RCG [rcg] focuses on unconditional generation, which first use the features from a self-supervised image encoder to serve as the ‘label’ for image generation by a second generator. Diff-AE [diff-ae] uses a learnable encoder for discovering the high-level semantics and a diffusion model for modeling stochastic; this dual-encoding improves the realism of the generated images on attribute manipulation and image interpolation tasks. Different from these works, our study focuses on the class-conditional diffusion transformer’s pre-training and exploiting representation guidance in itself and its own training paradigm. But we also believe our plug-and-play method can be easily applied to other tasks with benefits.
@@
Knowledge distillation for diffusion model. Knowledge distillation [kd] is also widely used in diffusion models. Its purposes can roughly be divided into two types: improving generation performance and accelerating inference sampling. To achieve the first goal, A more powerful pretrained model is often used to guide the diffusion model during training [flux1-lite, repa, tinyfusion]. For example, TinyFusion [tinyfusion] combines progressive distillation with architecture-specific optimizations for U-Net backbones, enabling deployment on edge devices. REPA [repa] distills the knowledge from a large-scale, pretrained representation foundation model [dinov2, clip] and finds that this distillation can improve both training efficiency and generation quality. Meanwhile, to achieve the second goal, the student and teacher models will be initialized by pre-trained models (here it can be one model or two different models), then the training target is to make the prediction of the student model with fewer steps conform to that of the teacher model with multiple steps [lcm, hypersd, guided-distill, cm, dmdr], thereby achieving the result of accelerated sampling and generation speed. For example, CM [cm] constrains the student’s outputs of adjacent points on a sampling path to be consistent with teacher’s. LCM [lcm] leverages this idea in the latent space. Our study also have some similarities, while we do not need a pretrained model and improve the generation performance of diffusion transformers by applying proposed SRA.
Appendix L More Qualitative Results (CFG 4.0)
@@
Below we show some uncurated generation results on ImageNet $\times256$ from the SiT-XL + SRA. We use classifier-free guidance with $w$ = 4.0.